OpenStudy (anonymous):

Evaluate cot 45° without using a calculator by using ratios in a reference triangle.

OpenStudy (jdoe0001):

hint: \(\bf cot(\theta)=\cfrac{adjacent}{opposite}=\cfrac{y}{x}=\cfrac{{\color{blue}{ cos}}(\theta)}{{\color{brown}{ sin}}(\theta)}\qquad thus \\ \quad \\ cot(45^o)=\cfrac{adjacent}{opposite}=\cfrac{y}{x}=\cfrac{{\color{blue}{ cos}}(45^o)}{{\color{brown}{ sin}}(45^o)}\)

OpenStudy (anonymous):

thanks jdoe0001

OpenStudy (jdoe0001):

yw